Without parentheses, the expression 8 + 30 ÷ 2 + 4
equals 27. where would you Place parentheses in the expression
so that it equals 13; then 23

Answer by josgarithmetic(39620)

23 or 13 ?...

(8 + 30) ÷ (2 + 4)=38/6=19/3
(8 + 30 ÷ 2) + 4 =23+4=27
8 + (30 ÷ 2 + 4)= 8+19=27

Can you find other parentheses arrangements?

8 + 30 ÷ (2 + 4) =8+30/6=8+5=13 THIS IS ONE OF YOUR NUMBERS.

(8 + 30) ÷ 2 + 4=38/2+4=19+4=23 THIS IS ALSO ONE OF YOUR NUMBERS.